OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4?

OpenStudy (anonymous):

what are your options to choose from?

OpenStudy (anonymous):

y = - 3 x -4

OpenStudy (arabpride):

y = mx + b so you equation would look some thing like: y=-3x - 4

OpenStudy (anonymous):

a.y = –3x – 4 b.y = 3x + 4 c.y = 3x – 4 d.y = –3x + 4